OpenStudy (klimenkov):

Compute \[\huge{\sqrt{2}^{\sqrt{2}^{\sqrt{2}^{\sqrt{2}^{\sqrt{2}^\ldots}}}}}\]

OpenStudy (anonymous):

I'm not sure your expression makes any sense. Is the continued exponentiation finite or not?

OpenStudy (klimenkov):

It is infinite.

OpenStudy (anonymous):

Ah, thanks for fixing it

OpenStudy (klimenkov):

Hope, you like it now. What will it be?

OpenStudy (anonymous):

faliure

OpenStudy (anonymous):

I'm pretty sure it's infinite. \[\left(\left(\left(\sqrt2^\sqrt2\right)^\sqrt2\right)^\sqrt2\right)^{\ldots}\\ \left(\left(\left( \left(2^\frac{1}{2}\right) ^\sqrt2\right)^\sqrt2\right)^\sqrt2\right)^{\ldots}\\ \left(\left( \left(2^\frac{\sqrt2}{2}\right) ^\sqrt2\right)^\sqrt2\right)^{\ldots}\\ \left( \left(2^\frac{2}{2}\right)^\sqrt2\right)^{\ldots}\\ \left(2^\sqrt2\right)^{\ldots}\\ \left(\left(2^\sqrt2\right)^\sqrt2\right)^{\ldots}\\ \left(\left(\left(2^2\right)^\sqrt2\right)^\sqrt2\right)^{\ldots}\\ \left(\left(4^\sqrt2\right)^\sqrt2\right)^{\ldots}\\ \left(4^2\right)^{\ldots}\] Continuing in this way, the base keeps getting squared.

OpenStudy (klimenkov):

The order is wrong. Compare: \[2^{2^3}=2^8=256,\quad (2^2)^3=4^3=64.\]

OpenStudy (shubhamsrg):

why don't let whole of it = x ?

OpenStudy (shubhamsrg):

then if, you observe properly, you'll be reduced to [ sqrt(2) ]^x = x

OpenStudy (klimenkov):

Nice, what will it be?

OpenStudy (shubhamsrg):

hmm, we get stuck ther! -_-

OpenStudy (shubhamsrg):

there*

OpenStudy (shubhamsrg):

2^(x/2) = x => 2^x = x^2 symmetrically, there is a solution , see that ?

hartnn (hartnn):

integer solutions, 2 and 4...

OpenStudy (shubhamsrg):

which one will it be ?

OpenStudy (klimenkov):

Now tell me, what is the answer: 2 or 4?

hartnn (hartnn):

i think 2

OpenStudy (klimenkov):

Why not 4?

hartnn (hartnn):

because while squaring, we attached an extra root, x=4

OpenStudy (klimenkov):

Hm... \(\sqrt 2^4=4\). Why do you call it "extra root"?

hartnn (hartnn):

:O .. let me think more...

OpenStudy (shubhamsrg):

2 also fits in though.

OpenStudy (shubhamsrg):

on the calculator, it approaches 2, I am not too sure how can we prove that ?

Parth (parthkohli):

Let \(\epsilon > 0\)

OpenStudy (klimenkov):

A medal will be given only if you prove this.

OpenStudy (shubhamsrg):

-_-

OpenStudy (shubhamsrg):

arey to hell with the medal !

hartnn (hartnn):

lol

Parth (parthkohli):

LOL!!

OpenStudy (klimenkov):

Can somebody say what means "arey to hell" ?

OpenStudy (shubhamsrg):

ignore "arey" I meant To Hell with the medal.

OpenStudy (klimenkov):

Ok. I found an interesting proof to this. Hope, you will find another interesting proof.

OpenStudy (shubhamsrg):

unable to prove, what proof you got ?

OpenStudy (shubhamsrg):

I mean I am unable to prove*

OpenStudy (klimenkov):

 Hope, you understand.